# Good News for Polynomial Root-finding

### v35

@@ 12.1 Simple reduction to a single e/i test in the case of an isolated disc
12.2 $m$-test by means of e/i tests for any disc
Our alternative reduction of $m$-test for $1\leq m\leq d$ to a bounded number of e/i tests can be applied in the case of any disc, not necessarily isolated. If at least one of the e/i tests fails, then definitely there is a root in a narrow annulus about the boundary circle of an input disc, and hence the internal disc of the annulus would not contain $m$ roots. Otherwise we would obtain a root-free concentric annulus about the boundary circle of an input disc and then reduce $m$-test to root-counting in an isolated disc, which we can perform at a low cost (see, e.g., claim (ii) of Prop. 3.1).
Algorithm 12.1.
Reduction of $m$-test to e/i tests.
INPUT: a polynomial $p(x)$ of a degree $d$, two values $\bar{\theta}>1$ and $\bar{\sigma}>1$, an integer $m$ such that $1\leq m\leq d$ and #$\#(D(0,\bar{\sigma}))\leq m$, and two algorithms $\mathbb{ALG}~{}0$ and $\mathbb{ALG}~{}1$, which a black box polynomial perform $\bar{\sigma}$-soft e/i test for in any disc and root-counting in an isolated disc, respectively.
INITIALIZATION (see Fig. 6): Fix a positive $\rho$ and compute the integer
$v=\lceil 2\pi/\phi\rceil$ (12.4)
where $\phi$ is the angle defined by an arc of $C(0,1)$ with end points at the distance $\rho$. Then compute the values
$c_{j}=\exp(\phi_{j}{\bf i}),~{}j=0,1,\dots,v,$ (12.5)
$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c_{5}$$c_{6}$$c_{7}$$c_{8}$$c_{9}$$c_{0}$
Figure 6: $v=9$, $|c_{j}-c_{j-1}|=\rho>|c_{9}-c_{0}|$ for $j=1,2,\dots,9$.
To prove correctness of Alg. 12.1 we need the following result.
Theorem 12.1.
Proof.
$c_{0}$$c_{1}$$z$$z^{\prime}$
Corollary 12.1.
Correctness of Alg. 12.1. (i) If algorithm $\mathbb{ALG}~{}1$ outputs inclusion in at least one of its applications, then one of the discs $D(c_{j},\bar{\sigma}\rho)$ contains a root, and then Alg. 12.1 correctly reports that $r_{d-m+1}\geq 1-\sigma\bar{\rho}$.
(ii) Otherwise Cor. 12.1 holds, and then correctness of Alg. 12.1 follows from correctness of $\mathbb{ALG}~{}1$.
Computational complexity. Alg. 12.1 invokes algorithm $\mathbb{ALG}~{}0$ at most $v+1$ times for $v$ of (12.4). If $\mathbb{ALG}~{}0$ outputs exclusion $v+1$ times, then algorithm $\mathbb{ALG}~{}1$ is invoked.
Remark 12.1.
One can modify Alg. 12.1 as follows (see [60]): instead of fixing a positive $\rho$ and then computing $v$ of (12.4) fix an integer $v>0$, apply algorithm $\mathbb{ALG}~{}0$ at $v$ equally spaced points $c_{j}=\exp(\frac{2j\pi{\bf i}}{v}),~{}j=0,1,\dots,v-1$, let $\phi=\frac{2\pi}{v}$, and then define $\rho$. We also arrive at this algorithm if we compute or guess this value $\rho$ and then apply Alg. 12.1, but notice that $c_{v}=c_{0}$ for $c_{j}$ of (12.5) and only invoke algorithm $\mathbb{ALG}~{}0$ for $c_{0},\dots,c_{v-1}$.
@@ 13.1 Doubling computational precision towards its a posteriori estimation
Given an algorithm $\mathbb{A}$ for black box polynomial computations and target tolerance TOL to output error, how should one choose computational precision? A simple practical recipe of [10, 14, 21] was the basis for highly successful empirical performance of MPSolve: namely, first apply the algorithm $\mathbb{A}$ with a lower precision and then recursively double it until an output error decreases below TOL; at that point it outputs precision within a factor of two from optimal, thus also supporting near optimal a posteriori bit operation complexity estimate for algorithm $\mathbb{A}$.
@@ 13.2 A priori estimates for computational precision
In the rest of this section we deduce a priori estimate for computational precision by proving that a precision of $O(b\log(d))$ bits is sufficient for supporting correctness of our subdivision root-finders. Such an estimate is deduced in [73] based on advanced non-trivial techniques, but we only use a very small part of them because the paper [73] works with higher order divided differences, while we only use the first order ones combined with Las Vegas randomization and some rather routine techniques for numerical error estimation.
Estimating computational precision of our soft e/i tests we assume that they are applied to the unit disc $D(0,1)$ because we can extend the results to any disc $D(c,\rho)$ by applying map (1.4) and increasing the upper bounds on the precision by a factor of $\log_{2}(1/\rho)$, which is at most $b$ because we only need to handle discs of radii $\rho>\epsilon=1/2^{b}$ and not $\epsilon$-discs. Therefore, our bound $O(\log(d))$ on the computational precision of our e/i tests for the disc $D(0,1)$ will imply the target bound $O(b\log(d))$ for testing any disc.
We also make Assumption 0 (Relative error bound) that we can request and obtain from evaluation oracle (black box subroutine) the values of $p$ with a relative error bound in $1/d^{\nu}$ for any fixed constant $\nu>0$ of our choice; we can represent these value with the bit-precision $\log_{2}(d)+\log_{2}(\nu)$.
@@
Multiplication with a precision $s$ contributes a relative error at most $2^{s}$ (cf. [20, Eqn. A.3 of Ch. 3]) and otherwise just sums the relative error bounds of the factors because we can ignore the dominated impact of higher order errors. Hence we deduce that $h=O(\log(d))$ multiplications for root-lifting to their $h$th powers increase these bounds at most by a factor of $h$, which is immaterial since we allow relative errors of order $1/d^{\nu}$ for a constant $\nu>0$ of our choice.
We only need to apply this analysis to approximation of NIR$(x)$ for $p_{h}(x)$ based on expression (1.6).
In our analysis we also use the following
Assumption 1 (Isolation Assumption): Whp all $q$ evaluation points $x\in C(0,1)$ of an e/i test for the unit disc $D(0,1)$ lie at least at the distance $\phi:=\frac{2\pi}{mqw}$ from all zeros of $p$ for any fixed $w>1$ of our choice.
We justify this assumption by applying our e/i tests under random rotation $x\mapsto vx$ for the random variable $v$ specified in our next proposition.
@@
Let $\phi_{g}:=\min_{j=1}^{d}|v\zeta^{g}-z_{j}|$ and let $\phi:=\min_{g=0}^{q-1}\phi_{g}$ for the zeros $z_{j}$ of $p$, $\zeta$ of (1.10) and random variable $v$ sampled from a fixed arc of the circle $C(0,1)$ of length $2\pi/q$ under the uniform probability distribution on that arc. Let $i(D(0,1))=m$. Then $\phi\leq\frac{2\pi}{mqw}$ for $w>1$ with a probability at most $\frac{\pi}{w}$.
@@
Next we prove that precision of $O(\log(d))$ bits supports our e/i tests under Assumptions 0 and 1 by applying Prop. 13.1 for sufficiently large $w=d^{O(1)}$. We will (i) prove that $|{\rm NIR}(z)|=d^{O(1)}$, (ii) approximate this value within $d^{\nu}\delta$ for any fixed constant $\nu$ and for $\delta$ of (1.6) chosen such that $\log(\frac{1}{\delta})=O(\log(d))$, and (iii) estimate that such a bound withstands the impact of rounding errors of computations for (1.6).
@@
In the proof of the next theorem we use the following lemma, which is [73, Fact 3.5].
@@
Under the assumptions of Prop. 13.2 it holds that
@@
The claimed bound on $|\xi^{\prime}|$ follows from (1.13). It remains to apply the first five lines of the proof of [73, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, respectively. Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
@@
Suppose that the values $t(y)$ have been computed by a black box oracle within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor. 13.2. Then one can ensure that $\beta<\frac{1}{d^{\bar{\eta}}}$ for any fixed constant $\bar{\eta}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $O(\log(d))$ bits, which we can do under Assumption 0.
@@
By extending Cor. 13.2 obtain $|\alpha^{\prime}|\leq\frac{1}{d^{O(1)}}$ and deduce from (1.13) and Assumptions 0 and 1 that $|\frac{t^{\prime}(x)}{t(x)}|\leq\frac{1}{d^{O(1)}}$ (cf. the proof of Prop. 6.1). Hence $w:=|\frac{t(x-\delta)}{t(x)}|\leq d^{O(1)}$ and $|\beta\delta|\leq|w|\frac{|\nabla(x)|+|\nabla(x-\delta)|}{1-|\nabla(x)|}$.
@@
Combine Cors. 13.1 and 13.2 and Prop. 13.4 to obtain
@@
We derived our precision bound $O(\log(d))$ (in the case of root-finding in the unit disc $D(0,1)$) based on operating entirely with NIR$(x)$. For comparison, if we relied on computing $p(x)$ and on the bound $\min_{x:~{}|x|=1}(|p(x)|/\sum_{i=0}^{d}|p_{i}|)\geq(\frac{\theta-1}{2\theta})^{d}$ of [130, Eqn. (9.4)], we would have only obtained a bound of order $m^{2}d$ on the computational precision of our root-finders.
@@ 14 Approximation of matrix eigenvalues
Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [134, Thm. 1.3.2]: $|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|$ for some $i$, $1\leq i\leq d$.
@@
Our root-finders are essentially reduced to evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$. Hence we can rely on the following expressions for approximation of the eigenvalues of a matrix or a matrix polynomial:
@@
where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [19, Eqn. (5)]. [^36] These expressions have been obtained based on Jacobi’s formula ${\rm d}(\ln(\det(A)))={\rm trace}(A^{-1}{\rm d}A),$ which links differentials of a nonsingular matrix $A$ and of its determinant. Based on the above expressions one can reduce approximation of eigenvalues of a matrix or a matrix polynomial to its inversion. This computation tends to be more numerically stable than computing determinants and is significantly accelerated the highly important case where an input matrix or matrix polynomial can be inverted fast, e.g., is data sparse.
@@ 15 Bit operation complexity of general polynomial root-finding
We could simply combine the precision bound $O(b\log(d))$ of Sec. 13 for our e/i tests with the estimate $O(bm)$ of Cor. 11.2 for the number of evaluation points for our root-finders, the bound of $2d$ ops for the evaluation of $p(x)$, and (1.8) for bit-operation cost of an op and then would arrive at the bit-operation complexity bound $\tilde{O}(b^{2}dm)$ for our root-finders. [^37] Recall that under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $m$.
Next, however, we decrease this bound roughly by a factor of $d$ by using benefits of fast multipoint evaluation of a polynomial by means of the algorithm of Fiduccia [40] and Moenck and Borodin [81] (see [18, Sec. 4.5] for historical account), non-trivially extended in [78] to support the following theorem:
@@
For $q=d$ this is [78, Thm. 2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed by Guillaume Moroz in [78]. Trivially extend this theorem to the cases where $q>d$ (by considering $t(x)$ a polynomial of degree $q$ with $q-d$ leading coefficients 0) and where $q<d$ (by removing $d-q$ points $x_{i}$, with no increase of the complexity bound). ∎
@@
due to [65, Thm. 3.9, Alg. 5.3, and Appendix A.3] by Peter Kirrinnis, which relied on combining his and Schönhage’s study of bit operation complexity of basic polynomial computations (cf. [130]) with the algorithm of [40, 81]. [^38] Precision of computing required in the latter algorithm grows fast as $d$ increases; e.g., the IEEE standard double precision of the customary numerical computations is not enough to support that algorithm already for $d>50$; for $b=O(\log(d))$ one can fix this deficiency by applying the Fast Multipole Method of Greengard and Rokhlin [48, 27, 101, 102] to the equivalent task of multiplication of Vandermonde matrix by a vector. One can, however, support bounds of Eqns. (15.1) and (15.2) on the overall bit operation complexity by computing with multiple precision for larger values $b$. Moroz obtained [78, Thm. 2] by applying bound (15.2) to his lower degree polynomials approximating $p(x)$.
@@ 15.1 Reduction of our root-finders to multipoint polynomial approximation
Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ zeros of $p$, we approximate all these zeros within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with e/i tests of Sec. 11. Then every subdivision step is reduced to application of our e/i tests to $\bar{m}=O(m)$ discs $D(c_{\lambda},\rho_{\lambda})$, where $\rho_{\lambda}\geq R/2^{b}$ for all $\lambda$. Such a test amounts to approximation of $p(x)$ at $q$ equally spaced points $x$ on each circle $C(c_{\lambda},\rho_{\lambda})$.
In Thm. 15.1 a polynomial $t(x)$ with $||t(x)||_{1}\leq 1$ is approximated at some points $x\in D(0,1)$. To ensure the assumption about $x$, we scale the variable $x\mapsto Rx$ to map the discs $D(c_{\lambda},\rho_{\lambda}\sigma)\mapsto D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})\subseteq D(0,1)$, for $\bar{c}_{\lambda}=c_{\lambda}/R$, $\bar{\rho}_{\lambda}=\rho_{\lambda}/R\geq 1/2^{b}$, and all $\lambda$.
Then we write $t(x):=p(Rx)/\psi$ for $\psi=O(R^{d})$ such that $||t(x)||_{1}=1$ and approximate $t(x)$ within $1/2^{\bar{b}}$ or equivalently within the relative error bound $1/(|t(x)|2^{\bar{b}})$ at $\bar{m}$ points $x$ in $\bar{m}$ e/i tests. [^39] Under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $\bar{m}/\log(\bar{m})$ and then extend this decrease to the related subsequent estimates. Due to Cor. 13.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we achieve this by choosing $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
In our e/i tests we only need to approximate $t(x)$ at the points $x$ lying on $\theta_{\lambda}$-isolated circles where $\frac{1}{\theta_{\lambda}-1}=O(m)$ (cf. Assumption 1 of Sec. 13.2). Then $\log(\frac{1}{|t(x)|})=O(d\log(m))$ by virtue of the bound of [130, Eqn. (9.4)]). [^40] That bound follows from the estimate $|f|\geq 1/2^{d}$ of [130, Thm. 4.2], which holds provided that $t(x)=f\prod_{j=1}^{m}(x-x_{j})\prod_{j=m+1}^{d}(\frac{x}{x_{j}}-1)$, $||t(x)||_{1}=1$, $|x_{j}|<1$ for $j\leq m$, and $|x_{j}|>1$ for $j>m$.
@@
For $m=d$, the above bound on $\log(\frac{1}{|t(x)|})$ is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$. For $m<d$, [130, Thm. 4.5] implies a little stronger bounds, namely, $\log(\frac{1}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.

### v36

@@ 16.1 Doubling computational precision towards its a posteriori estimation
Given an algorithm $\mathbb{A}$ for black box polynomial computations and target tolerance TOL to the output errors, how should one choose computational precision? Here is a simple practical recipe from [10, 15, 22] and MPSolve for obtaining a posteriori estimate within a factor of two from optimal: first apply algorithm $\mathbb{A}$ with a lower precision, then recursively double it and stop where an output error decreases below TOL.
@@ 16.2 A priori estimates for computational precision
In the rest of this section we deduce a priori estimate of $O(b\log(d))$ bits for computational precision, proving that this precision supports correctness of our subdivision root-finders. Such an estimate is deduced in [76] based on advanced non-trivial techniques, of which we only use a very small part.
Estimating computational precision of our soft e/i tests we assume that they are applied to the unit disc $D(0,1)$ because we can extend the results to any disc $D(c,\rho)$ by applying map (1.3) and increasing the upper bounds on the precision by a factor of $\log_{2}(1/\rho)$. This factor is at most $b$ because we only handle discs of radii $\rho>\epsilon=1/2^{b}$ until we stop reaching $\epsilon$-discs. Therefore, our bound $O(\log(d))$ on the computational precision of our e/i tests for the disc $D(0,1)$ will imply the bound $O(b\log(d))$ for any disc.
We also make Assumption 0 that we can request and obtain from evaluation oracle (black box subroutine) the values of $p$ with a relative error bound in $1/d^{\nu}$ for any fixed constant $\nu>0$ of our choice; we can represent these values with the bit-precision $\log_{2}(d)+\log_{2}(\nu)$.
@@
Multiplication with a precision $s$ contributes a relative error at most $2^{s}$ (cf. [21, Eqn. A.3 of Ch. 3]) and otherwise just sums the relative error bounds of the factors ignoring the dominated impact of higher order errors. Hence we deduce that $h=O(\log(d))$ multiplications for root-lifting to their $h$th powers increase these bounds by at most a factor of $h$, which is immaterial since we allow relative errors of order $1/d^{\nu}$ for a constant $\nu>0$ of our choice.
We only apply this analysis to approximation of NIR$(x)$ for $p_{h}(x)$ based on (3.2) and also use the following
Assumption 1: Whp all $q$ evaluation points $x\in C(0,1)$ of an e/i test for the unit disc $D(0,1)$ lie at the distance at least $\phi:=\frac{2\pi}{mqw}$ from all roots for any fixed $w>1$ of our choice.
This assumption holds because we apply our e/i tests under random rotation $x\mapsto vx$ for the random variable $v$ specified in our next proposition.
@@
Let $\phi_{g}:=\min_{j=1}^{d}|v\zeta^{g}-z_{j}|$ and let $\phi:=\min_{g=0}^{q-1}\phi_{g}$ for the roots $z_{j}$, $\zeta$ of (3.5) and random variable $v$ sampled from a fixed arc of the circle $C(0,1)$ of length $2\pi/q$ under the uniform probability distribution on that arc. Let $i(D(0,1))=m$. Then $\phi\leq\frac{2\pi}{mqw}$ for $w>1$ with a probability at most $\frac{\pi}{w}$.
@@
Next we prove that a precision of $O(\log(d))$ bits supports our e/i tests under Assumptions 0 and 1 by applying Prop. 16.1 for sufficiently large $w=d^{O(1)}$. We will (i) prove that $|{\rm NIR}(z)|=d^{O(1)}$, (ii) approximate this value within $d^{\nu}\delta$ for any fixed constant $\nu$ and for $\delta$ of (3.2) chosen such that $\log(\frac{1}{\delta})=O(\log(d))$, and (iii) estimate that such a bound withstands the impact of rounding errors of computations for (3.2).
@@
In the proof of the next theorem we use the following lemma, which is [76, Fact 3.5].
@@
Under the assumptions of Prop. 16.2 it holds that
@@
The claimed bound on $|\xi^{\prime}|$ follows from (3.1). It remains to apply the first five lines of the proof of [76, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, respectively. Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
@@
Suppose that the values $t(y)$ have been computed by a black box oracle within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor. 16.2. Then one can ensure that $\beta<\frac{1}{d^{\bar{\eta}}}$ for any fixed constant $\bar{\eta}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $O(\log(d))$ bits, which we can do under Assumption 0.
@@
By extending Cor. 16.2 obtain $|\alpha^{\prime}|\leq\frac{1}{d^{O(1)}}$ and deduce from (3.1) and Assumptions 0 and 1 that $|\frac{t^{\prime}(x)}{t(x)}|\leq\frac{1}{d^{O(1)}}$ (cf. the proof of Prop. 10.1). Hence $w:=|\frac{t(x-\delta)}{t(x)}|\leq d^{O(1)}$ and $|\beta\delta|\leq|w|\frac{|\nabla(x)|+|\nabla(x-\delta)|}{1-|\nabla(x)|}$.
@@
Combine Cors. 16.1 and 16.2 and Prop. 16.4 to obtain
@@
We derived our precision bound $O(\log(d))$ (in the case of root-finding in the unit disc $D(0,1)$) based on operating entirely with NIR$(x)$. For comparison, if we relied on computing $p(x)$ and on the bound $\min_{x:~{}|x|=1}(|p(x)|/\sum_{i=0}^{d}|p_{i}|)\geq(\frac{\theta-1}{2\theta})^{d}$ of [134, Eqn. (9.4)], we would have only obtained a bound of order $m^{2}d$ on the computational precision of our root-finders.
@@ 17 Approximation of matrix eigenvalues
Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [138, Thm. 1.3.2]: $|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|$ for some $i$, $1\leq i\leq d$.
@@
Our root-finders are essentially reduced to the evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$, and we can rely on the following expressions for approximation of the eigenvalues of a matrix or a matrix polynomial:
@@
where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [20, Eqn. (5)]. [^28] These expressions have been obtained based on Jacobi’s formula ${\rm d}(\ln(\det(A)))={\rm trace}(A^{-1}{\rm d}A),$ which links differentials of a nonsingular matrix $A$ and of its determinant. Based on the above expressions one can reduce approximation of eigenvalues of a matrix or a matrix polynomial to its inversion. This computation tends to be more numerically stable than computing determinants and is significantly accelerated the highly important case where an input matrix or matrix polynomial can be inverted fast, e.g., is data sparse.
@@ 18 Bit operation complexity of general polynomial root-finding
We could simply combine the precision bound $O(b\log(d))$ of Sec. 16 for our e/i tests with the estimate $O(bm)$ of Cor. 14.2 for the number of evaluation points for our root-finders, the bound of $2d$ ops for the evaluation of $p(x)$, and (12.5) for bit-operation cost of an op and then would arrive at the bit-operation complexity bound $\tilde{O}(b^{2}dm)$ for our root-finders. [^29] Recall that under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $m$.
Next, however, we decrease this bound roughly by a factor of $d$ by using benefits of fast multipoint evaluation of a polynomial by means of the algorithm of Fiduccia [42] and Moenck and Borodin [84] (see [19, Sec. 4.5] for historical account), non-trivially extended in [81] to support the following theorem:
@@
For $q=d$ this is [81, Thm. 2], based on using novel data structure for hyperbolic approximation of $p$, proposed and analyzed by Guillaume Moroz in [81]. Trivially extend this theorem to the cases where $q>d$ (by considering $t(x)$ a polynomial of degree $q$ with $q-d$ leading coefficients 0) and where $q<d$ (by removing $d-q$ points $x_{i}$, with no increase of the complexity bound). ∎
@@
due to [68, Thm. 3.9, Alg. 5.3, and Appendix A.3] by Peter Kirrinnis, which relied on combining his and Schönhage’s study of bit operation complexity of basic polynomial computations (cf. [134]) with the algorithm of [42, 84]. [^30] Precision of computing required in the latter algorithm grows fast as $d$ increases; e.g., the IEEE standard double precision of the customary numerical computations is not enough to support that algorithm already for $d>50$; for $b=O(\log(d))$ one can fix this deficiency by applying the Fast Multipole Method of Greengard and Rokhlin [50, 28, 105, 106] to the equivalent task of multiplication of Vandermonde matrix by a vector. One can, however, support bounds of Eqns. (18.1) and (18.2) on the overall bit operation complexity by computing with multiple precision for larger values $b$. Moroz obtained [81, Thm. 2] by applying bound (18.2) to his lower degree polynomials approximating $p(x)$.
@@ 18.1 Reduction of our root-finders to multipoint polynomial approximation
Given a complex $c$, a pair of positive $b$ and $\rho$, and $d+1$ coefficients of a polynomial $p=p(x)$ of (1.1) such that the disc $D(c,\rho)$ is isolated and contains precisely $m$ roots, we approximate all these roots within $R/2^{b}$ for $R=|c|+\rho$ by applying subdivision root-finding iterations with e/i tests of Sec. 14. Then every subdivision step is reduced to application of our e/i tests to $\bar{m}=O(m)$ discs $D(c_{\lambda},\rho_{\lambda})$, where $\rho_{\lambda}\geq R/2^{b}$ for all $\lambda$. Such a test amounts to approximation of $p(x)$ at $q$ equally spaced points $x$ on each circle $C(c_{\lambda},\rho_{\lambda})$.
In Thm. 18.1 a polynomial $t(x)$ with $||t(x)||_{1}\leq 1$ is approximated at some points $x\in D(0,1)$. To ensure the assumption about $x$, we scale the variable $x\mapsto Rx$ to map the discs $D(c_{\lambda},\rho_{\lambda}\sigma)\mapsto D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})\subseteq D(0,1)$, for $\bar{c}_{\lambda}=c_{\lambda}/R$, $\bar{\rho}_{\lambda}=\rho_{\lambda}/R\geq 1/2^{b}$, and all $\lambda$.
Then we write $t(x):=p(Rx)/\psi$ for $\psi=O(R^{d})$ such that $||t(x)||_{1}=1$ and approximate $t(x)$ within $1/2^{\bar{b}}$ or equivalently within the relative error bound $1/(|t(x)|2^{\bar{b}})$ at $\bar{m}$ points $x$ in $\bar{m}$ e/i tests. [^31] Under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $\bar{m}/\log(\bar{m})$ and then extend this decrease to the related subsequent estimates. Due to Cor. 16.3 it is sufficient to ensure the bound $1/(|t(x)|2^{\bar{b}})=1/d^{O(1)}$; we achieve this by choosing $\bar{b}\geq\log(\frac{1}{|t(x)|})+O(\log(d))$.
In our e/i tests we only need to approximate $t(x)$ at the points $x$ lying on $\theta_{\lambda}$-isolated circles where $\frac{1}{\theta_{\lambda}-1}=O(m)$ (cf. Assumption 1 of Sec. 16.2). Then $\log(\frac{1}{|t(x)|})=O(d\log(m))$ by virtue of the bound of [134, Eqn. (9.4)]). [^32] That bound follows from the estimate $|f|\geq 1/2^{d}$ of [134, Thm. 4.2], which holds provided that $t(x)=f\prod_{j=1}^{m}(x-x_{j})\prod_{j=m+1}^{d}(\frac{x}{x_{j}}-1)$, $||t(x)||_{1}=1$, $|x_{j}|<1$ for $j\leq m$, and $|x_{j}|>1$ for $j>m$.
@@
For $m=d$, the above bound on $\log(\frac{1}{|t(x)|})$ is sharp up to a constant factor and is reached for the polynomial $t(x)=(x+1-\frac{1}{m})^{d}/(2-\frac{1}{m})^{d}$, $\rho_{\lambda}=1$, and $x=-1$. For $m<d$, [134, Thm. 4.5] implies a little stronger bounds, namely, $\log(\frac{1}{|t(x)|})=O(d+m\log(m))$ and hence $\bar{b}=O(m\log(m)+d+b)$ because all discs $D(\bar{c}_{\lambda},\bar{\rho}_{\lambda})$ lie in $\theta$-isolated unit disc $D(0,1)$ for $\theta-1$ exceeding a positive constant.